### v1

# Efficient Retrieval with Learned Similarities

Abstract
Retrieval plays a fundamental role in recommendation systems, search, and natural language processing by efficiently finding relevant items from a large corpus given a query. Dot products have been widely used as the similarity function in such retrieval tasks, thanks to Maximum Inner Product Search (MIPS) that enabled efficient retrieval based on dot products. However, state-of-the-art retrieval algorithms have migrated to learned similarities. Such algorithms vary in form; the queries can be represented with multiple embeddings, complex neural networks can be deployed, the item ids can be decoded directly from queries using beam search, and multiple approaches can be combined in hybrid solutions. Unfortunately, we lack efficient solutions for retrieval in these state-of-the-art setups. Our work investigates techniques for approximate nearest neighbor search with learned similarity functions. We first prove that Mixture-of-Logits (MoL) is a universal approximator, and can express all learned similarity functions. We next propose techniques to retrieve the approximate top $K$ results using MoL with a tight bound. We finally compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
1 Introduction
Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors. Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items (Covington et al., 2016; Gillick et al., 2018; Li et al., 2019; Lewis et al., 2020; Borgeaud et al., 2022). In many concrete use cases, such as vector databases (Johnson et al., 2021), the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
Despite the popularity of dot products and numerous work done to improve their efficiency (Jegou et al., 2011; Shrivastava and Li, 2014; Malkov and Yashunin, 2020; Chern et al., 2022), state-of-the-art retrieval algorithms have long moved to various learned similarity functions. Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values (Li et al., 2019; Khattab and Zaharia, 2020). As another example, Probabilistic Label Trees (PLTs) (Jasinska et al., 2016) and Tree-based Deep Models (TDMs) (Zhu et al., 2018; Zhuo et al., 2020) map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf. More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups (Tay et al., 2022; Wang et al., 2022). Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products (Gao et al., 2021). Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms (He et al., 2017; Yang et al., 2018; Santhanam et al., 2022; Zhai et al., 2023).
Supporting efficient retrieval with these diverse learned similarities is challenging. Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps (Zhu et al., 2018; Zhai et al., 2023), while large scale recommendation systems commonly need to handle billions of items (Eksombatchai et al., 2018; Li et al., 2019) with low end-to-end latency. When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force (He et al., 2017) or heuristics (Zhai et al., 2023). What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface for these techniques makes it even more difficult for designing a general solution for efficient retrieval.
Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage. Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them. Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$). This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases (Johnson et al., 2021; Chern et al., 2022), and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data (Covington et al., 2016; Li et al., 2019; Gao et al., 2021; Zhai et al., 2024).
This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL. To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities. We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions. We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound. Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS (Jegou et al., 2011; Shrivastava and Li, 2014). Importantly, our approach efficiently utilizes modern accelerators like GPUs due to higher arithmetic intensity. We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
2 Expressiveness of Mixture of Logits
In this section, we describe Mixture of Logits (MoL) and show that MoL is expressive enough to represent any learned similarity function. Table 1 summarizes the notations in this paper.
We first describe the learned similarity function with Mixture of Logits (MoL).
Notation Description
$q$ ($Q$, $|Q|$) query (set of queries, number of queries)
$x$ ($X$, $|X|$) item (set of items, number of items)
$\phi(q,x)$ the learned similarity function, i.e., Mixture-of-Logits (MoL).
$P$ ($P_{q}$, $P_{x}$) number of component-level embeddings. With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of component-level embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
$\pi_{p}(q,x)$ ($\pi_{p_{q},p_{x}}(q,x)$) weight for the $p$-th (or $p_{q}$-th by $p_{x}$-th with outer product) embedding set for $(q,x)$.
$f(q)$ ($f_{p}(q)$) learned embedding for the query ($p$-th component-level query embedding)
$g(x)$ ($g_{p}(x)$) learned embedding for the item ($p$-th component-level item embedding)
$\langle f(q),g(x)\rangle$ the dot product similarity function; $\langle f(q),g(x)\rangle=g(x)^{T}f(q)$.
Table 1: Table of Notations.
Mixture of Logits (MoL)
MoL applies an adaptive gating weight $\pi_{p}(q,x)$ to $P$ groups of component-level dot products $\langle f_{p}(q),g_{p}(x)\rangle$. Following Zhai et al. (2023):
$\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ (1)
To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embeddings (where $P_{q}\times P_{x}=P$) and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
$\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$ (2)
We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper as the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
Figure 1: Mixture-of-logits (MoL) learned similarity.
Now we show that any high rank matrix can be decomposed into a mixture of logits with low rank matrices. Specifically, without loss of generality, we prove the following:
Theorem 1.
[…]
First, the theorem trivially holds if the rank of $A$ is less than or equal to $d$:
Lemma 1.
MoL decomposition when $Rank(A)\leq d$: Let $A$ be a matrix as defined in Theorem 1. If the rank of $A$ is less than or equal to $d$, then we have $A=\pi\circ A$, where $\pi(i,j)=1$ for any $1\leq i\leq n,1\leq j\leq m$.
Then we prove for the case where the rank of $A$ is greater than $d$. Without loss of generality, we prove the case where the matrix has full rank, i.e., $Rank(A)=n$:
Lemma 2.
MoL decomposition when $Rank(A)=n$: Let $A$ be a matrix as defined in Theorem 1. Then there exists $\pi,B_{1},B_{2}$ such that $|A-(\pi\circ B_{1}+(1-\pi)\circ B_{2})|<\epsilon$, where $Rank(B_{1})\leq d$, $Rank(B_{2})\leq d$, and $0\leq\pi(i,j)\leq 1$ for $1\leq i\leq n,1\leq j\leq m$.
Proof.
Because $A$ is a matrix of rank $n$, it can be rewritten as $A=UI_{n}V$, where $I_{n}$ is an identity matrix with rank $n$. Thus, $A_{ij}=\sum_{k=1}^{n}U_{ik}V_{kj},1\leq i\leq n,1\leq j\leq m$. Let $A^{\prime}$ be a matrix of $n\times m$, where $A^{\prime}_{ij}=\lambda_{ij}\cdot\sum_{k=1}^{d}U_{ik}V_{kj}$ for $1\leq i\leq n,1\leq j\leq m$. Here, $\lambda_{ij}=1+\frac{\sum_{k=d+1}^{n}U_{ik}V_{kj}}{\sum_{k=1}^{d}U_{ik}V_{kj}}$ if $\sum_{k=1}^{d}U_{ik}V_{kj}\neq 0$, otherwise $\lambda_{ij}=1+\frac{\sum_{k=d+1}^{n}U_{ik}V_{kj}}{\epsilon}$. Thus, we have $|A-A^{\prime}|\leq\epsilon$.
Let $\lambda_{min}=\min{\lambda_{ij}}$, and $\lambda_{max}=\max{\lambda_{ij}}$. Let $B_{1}=\lambda_{min}UI_{d}V$, $B_{2}=\lambda_{max}UI_{d}V$, where $I_{d}$ is the identity matrix of rank $d$. We have $A^{\prime}_{ij}=\lambda_{ij}\sum_{k=1}^{d}U_{ik}V_{kj}=\pi(i,j)\cdot B_{1ij}+(1-\pi(i,j))\cdot B_{2ij}$, where $\pi(i,j)=\frac{\lambda_{max}-\lambda_{ij}}{\lambda_{max}-\lambda_{min}}$. Because $\lambda_{min}\leq\lambda_{ij}\leq\lambda_{max}$, we have $0\leq\pi(i,j)\leq 1$.
Thus, we have constructed $B_{1},B_{2},\pi$ such that $|A-(\pi\circ B_{1}+(1-\pi)\circ B_{2})|=|A-A^{\prime}|\leq\epsilon$. ∎
Remark Here, we have shown that any high-rank matrix can be expressed as a mixture of logits of two low-rank matrices. Note that our decomposition is not intended to be used as a distillation of the original high-rank matrix. For instance, it is prohibitively expensive to populate the full matrix with a learned similarity function. In addition, the proof also does not indicate that having a sum of two mixture of logits is sufficient to train the embeddings and the learned similarity function. It is well-known that overparameterization is often necessary to enable efficient and performant training.
3 Retrieval Algorithms
In this section, we describe the problem of retrieving the top $K$ items with MoL and both exact and approximate algorithms for retrieval. Formally, we define the top $K$ retrieval problem as follows:
Definition 1.
Top $K$ with MoL: Let $q$ be a query and $X$ be a set of items, where both the query $q$ and each item $x\in X$ are associated with $P$ embeddings. Together we have $P$ pairs of embeddings, $(f_{p}(q),g_{p}(x)),1\leq p\leq P$. Let $\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ be the similarity score of $q,x$, where $x\in X$. The top $K$ query with MoL returns the $K$ items from $X$ with the highest $\phi(q,x)$s.
For approximate top $K$ retrieval with MoL, we define the gap of the approximate and exact top $K$ results as follows:
Definition 2.
Gap of approximate top $K$: Let $q$ be a query and $X_{K}$ be the set of exact top $K$ items for the query $q$ from a set of items $X$. Let $X^{*}$ be the approximate top $K$ results, where $X^{*}\subseteq X$. Let $S=\min\{\phi(q,x),x\in X^{*}\}$ and $S^{\prime}=\max\{\phi(q,x),x\in X_{K}\setminus X^{*}\}$. We call $S_{\Delta}=S^{\prime}-S$ the gap of the top $K$ with $X^{*}$.
3.1 Exact algorithm
The brute-force algorithm to retrieve the exact top $K$ with MoL is to evaluate $\phi(q,x)$ for each query $q$ and item $x$. This algorithm can be prohibitively expensive if the number of items is large. Instead, we describe a more efficient two-pass algorithm to retrieve the exact top $K$ items as shown in Algorithm 1.
Algorithm 1 Exact top $K$ algorithm.
1:query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
2:exact top $K$ items
3:$G\leftarrow\emptyset$
4:for $p\in P$ do
5:    $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
6:    $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p})$triangleright Retrieve top $K$ items for each pair of embeddings
7:$S_{min}\leftarrow\infty$
8:for $x\in G$ do
9:    $s\leftarrow MoL(q,x)$
10:    if $s<S_{min}$ then $S_{min}\leftarrow s$
11:$G^{\prime}\leftarrow\emptyset$
12:for $p\in P$ do
13:    $G^{\prime}\leftarrow G^{\prime}\cup RangeDotProduct(f_{p}(q),S_{min},X_{p})$$\triangleright$ Retrieve all items with a dot product higher than or equal to $S_{min}$ for $f_{p}(q)$ from $X_{p}$
14:return $BruteForceTopKMoL(q,G^{\prime})$$\triangleright$ Retrieve the top $K$ items from $G^{\prime}$ with MoL
We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4). Then we evaluate the MoL score of the items in the set $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8). Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11). Finally, we evaluate the MoL scores of the items in $G^{\prime}$ exhaustively, and return the top $K$ items with the highest scores (line 12).
We argue that Algorithm 1 retrieves the exact top $K$ items with MoL. Let $X_{K}$ be the set of the exact top $K$ items and $X^{\prime}$ be the result of Algorithm 1. Let $x\in X_{K}$ and $\phi(q,x)$ be the MoL score of $x$ and $q$. Since $x$ has the highest top $K$ score with MoL, $\phi(q,x)\geq S_{min}$. Since the MoL score is a weighted score over the dot product scores, we have $\max\{\langle f_{p}(q),g_{p}(x)\rangle,1\leq p\leq P\}\geq\phi(q,x)\geq S_{min}$. Since Algorithm 1 retrieves all the items with a dot product higher than or equal to $S_{min}$ of $q$ for each embedding $q_{p}$ (line 9-11), we have $x\in G^{\prime}$. Thus, $x\in X^{\prime}$. So we have shown that $X_{K}=X^{\prime}$.
3.2 Approximate algorithms
In the exact algorithm shown in Algorithm 1, we need to retrieve all the items with a dot product higher than or equal to a threshold. When the threshold is a loose filter of the item set, which may happen when the dot product scores are skewed, $G^{\prime}$ can be large, and the evaluation of MoL over a large set of candidate items can still be expensive. Here, we describe two heuristics to approximately retrieve the top $K$ items and analyze their gap against the exact top $K$ algorithm.
In both heuristics, we perform a two-stage retrieval as shown in Algorithm 2. In the first stage, we retrieve a set of $K^{\prime}$ candidate items that are potentially high in MoL score by using dot products (line 2). Note that $K^{\prime}$ can be larger than $K$, e.g., due to oversampling. In the second stage, we evaluate the MoL score of the candidate items and return the top $K$ items (line 3).
Algorithm 2 Approximate top $K$ algorithms.
1:a query $q$, a set of items $X$
2:approximate top $K$ items
3:function ApproxTopK($q,X,K,K^{\prime}$)
4:    $G\leftarrow TopKCandidate(q,X,K^{\prime})$triangleright Retrieve the top $K^{\prime}$ candidate items
5:    return $BruteForceTopKMoL(q,G,K)$$\triangleright$ Retrieve the top $K$ items with MoL
6:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
7:union of top $K$ items over $P$ component-level embeddings by dot product
8:function TopKPerEmbedding($q,X,K$)
9:    $G\leftarrow\emptyset$
10:    for $p\in P$ do
11:         $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
12:         $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p},K)$$\triangleright$ Retrieve the top $K$ items by dot product
13:    return $G$
14:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embedding $f_{p}(q),g_{p}(x)$
15:top $K$ items based on the averaged dot product, $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
16:function TopKAvg($q,X,K$)
17:    $q^{\prime}\leftarrow\sum_{p=1}^{P}f_{p}(q)$
18:    $X^{\prime}\leftarrow\{\sum_{p=1}^{P}g_{p}(x)/P,x\in X\}$triangleright Can be preprocessed and materialized.
19:    return $TopKDotProduct(q^{\prime},X^{\prime},K)$
Here, we describe two heuristics to retrieve the candidate item set:
Top $K$ per embedding.
Given a query $q$ and a set of items $X$, for each embedding set $p$, retrieve top $K$ items $X_{K,p}$ based on dot product ($\langle f_{p}(q),g_{p}(x)\rangle$). Return the union across $P$ queries.
The top $K$ per embedding heuristic returns the union of the top $K$ items for each embedding by dot product. We analyze the gap of this heuristic as follows:
Theorem 2.
Upper bound of the gap of top $K$ per embedding: Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S=\max\{\phi(q,x),x\in X_{K+1,p}\}$. Let $S_{min}$ be the top $K^{th}$ MoL score of the items in $\cup_{p}X_{K,p}$, then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
Remark Note that there exists an MoL such that $S_{\Delta}=S-S_{min}$, i.e., when $\pi_{p}(q,x)=1$ for $x_{p}=\operatorname*{arg\,max}_{x,p}\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p}\}$. Thus, the upper bound of $S_{\Delta}$ is tight.
Top $K$ average.
Given a query $q$ and a set of items $X$, return the top $K$ items with the highest average dot product $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
Note that the top $K$ average heuristic returns the exact top $K$ items when the gating weight distribution in MoL, $\pi$, is uniform.
This heuristic is interesting for two reasons. First, the items retrieved by this heuristic are likely to be the top $K$ items of MoL when the weight distribution is more balanced. This complements the heuristic that retrieves top $K$ per embedding. Second, in the setup where the set of embedding pairs is constructed as the outer product of the embeddings of a query and those of an item (Equation 2), the average dot product can be efficiently preprocessed and materialized for the items, and the computation of the top $K$ average is then agnostic to the number of embedding sets $P=P_{q}\times P_{x}$.
Formally, let $P=P_{q}\cdot P_{x}$ be the number of embedding pairs, where $P_{q}$ is the number of embeddings of a query $q$ and $P_{x}$ is that of an item $x$. The average dot product can be computed as
$\frac{1}{P}\cdot\sum_{p=1}^{P}{\langle f_{p}(q),g_{p}(x)\rangle}=\frac{1}{P}\cdot\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}{\langle f_{p_{q}}(q),g_{p_{x}}(x)\rangle}=\frac{1}{P}\cdot\left\langle\sum_{p_{q}=1}^{P_{q}}f_{p_{q}}(q),\sum_{p_{x}=1}^{P_{x}}g_{p_{x}}(x)\right\rangle$ (3)
Thus, we can preprocess the embeddings of the items and the query, so the number of embeddings accessed is $1$ per item for a given query, regardless of the overall number of component-level embeddings used by MoL, i.e., $P$.
Finally, we can combine the candidates retrieved from top $K$ per embedding group and the top $K$ average as the following:
Combined top $K$.
Given a query $q$, a set of items $X$, and $K$, return the union of the items from the top $K$ per embedding group across the $P$ groups and the top $K$ items from the top $K$ average.
Theorem 3.
Upper bound of the gap of combined top $K$. Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S_{min}$ as defined in Theorem 2. Let $X^{\prime}_{K}$ be the top $K$ items from top $K$ average. Let $S^{\prime}=\max\{\phi(q,x),x\in X\setminus(\cup_{p}X_{K,p}\cup X^{\prime}_{K})\}$. Then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
Remark Similar to Theorem 2, the upper bound of the gap is tight. In practice, we can configure the $K$ to be different for the two heuristics, i.e., $K_{1}$ and $K_{2}$. For example, when the weight distribution $\pi$ is more balanced, $K_{2}$ can be configured to be larger as the top $K$ average approach approximates MoL well while being more computationally efficient.
4 Evaluation
In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3. Our code and pretrained model checkpoints are publicly available at https://github.com/bailuding/rails.
4.1 Setup
Workloads.
We consider a classical retrieval task (next-item prediction) in recommendation systems (Covington et al., 2016; Zhu et al., 2018; Li et al., 2019; Zhai et al., 2023). We employ three widely used datasets, MovieLens (Harper and Konstan, 2015) (the 1M and 20M subsets), and Amazon Reviews (McAuley et al., 2015) (the largest Books subset). We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets (Kang and McAuley, 2018; Zhai et al., 2023). In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step. Table 2 shows the statistics of these three workloads.
Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ component-level embedding dimension
ML-1M 6,040 3,649 8 4 64
ML-20M 138,493 24,186 8 4 128
Books 694,897 674,044 8 8 32
Table 2: Workload statistics.
Model training.
We implemented MoL on top of two popular self-attention sequential encoder architectures, SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024). The models are implemented in PyTorch and are trained with 2x NVIDIA RTX 6000 Ada GPUs.
Method HR@1 HR@10 HR@50 HR@200 MRR
ML-1M SASRec .0610 .2806 .5437 .7498 .1307
SASRec + MoL .0672 .3019 .5711 .7650 .1431
HSTU .0774 .3306 .5897 .7832 .1571
HSTU + MoL .0875 .3412 .6013 .7877 .1692
ML-20M SASRec .0674 .2889 .5503 .7661 .1388
SASRec + MoL .0731 .3068 .5645 .7768 .1498
HSTU .0965 .3556 .6143 .8074 .1800
HSTU + MoL .1014 .3661 .6234 .8116 .1876
Books SASRec .0058 .0306 .0754 .1431 .0153
SASRec + MoL .0083 .0402 .0970 .1813 .0203
HSTU .0101 .0469 .1066 .1876 .0233
HSTU + MoL .0150 .0613 .1292 .2167 .0315
Table 3: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task.
Metrics.
We use Recall (Hit Rate) as the main metric. It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3). Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 21.4% in HR@1, an average of 13.7% in HR@10, and an average of 16.2% in MRR across the six scenarios, validating the importance of learned similarity functions. We evaluate our proposed retrieval algorithms relative to the best performing rows, specifically “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the following section.
4.2 Top $K$ retrieval performance
Method HR@1 HR@5 HR@10 HR@50 HR@100 Latency / ms
ML-1M BruteForce 1.00 1.00 1.00 1.00 1.00 .64$\pm$.03
TopKPerEmbd5 .838 .768 .681 .499 .437 .95$\pm$.03
TopKPerEmbd10 .932 .905 .842 .678 .618 .94$\pm$.05
TopKPerEmbd50 1.00 .997 .999 .977 .954 .95$\pm$.05
TopKPerEmbd100 1.00 1.00 1.00 .995 .990 1.24$\pm$.04
TopKAvg200 .980 .970 .967 .958 .948 .74$\pm$.06
TopKAvg500 .992 .986 .989 .990 .990 .71$\pm$.04
TopKAvg1000 .996 .994 .998 .997 .999 .72$\pm$.03
CombTopK5_200 1.00 .997 .988 .975 .960 1.09$\pm$.08
CombTopK50_500 1.00 .999 .999 .998 .997 1.13$\pm$.04
ML-20M BruteForce 1.00 1.00 1.00 1.00 1.00 3.08$\pm$.02
TopKPerEmbd5 .755 .702 .645 .467 .402 1.19$\pm$.05
TopKPerEmbd10 .946 .876 .817 .645 .564 1.17$\pm$.05
TopKPerEmbd50 .993 .986 .977 .931 .900 1.55$\pm$.03
TopKPerEmbd100 .998 .997 .994 .979 .966 2.26$\pm$.03
TopKAvg200 .992 .992 .992 .980 .958 .76$\pm$.04
TopKAvg500 .992 .993 .995 .995 .993 .77$\pm$.04
TopKAvg1000 .992 .994 .996 .998 .997 .75$\pm$.03
CombTopK5_200 1.00 .998 .998 .983 .962 1.32$\pm$.06
CombTopK50_500 1.00 1.00 1.00 .999 .996 1.85$\pm$.03
Books BruteForce 1.00 1.00 1.00 1.00 1.00 149.02$\pm$.49
TopKPerEmbd5 .907 .915 .809 .509 .396 20.17$\pm$.06
TopKPerEmbd10 .936 .986 .914 .698 .561 20.33$\pm$.06
TopKPerEmbd50 .993 .992 .994 .956 .902 21.09$\pm$.04
TopKPerEmbd100 1.00 .995 .996 .985 .959 22.53$\pm$.06
TopKAvg200 1.00 .978 .948 .845 .767 .68$\pm$.03
TopKAvg500 1.00 .992 .996 .919 .875 .68$\pm$.03
TopKAvg1000 1.00 1.00 .996 .963 .939 .84$\pm$.01
TopKAvg2000 1.00 1.00 1.00 .980 .968 1.04$\pm$.02
TopKAvg4000 1.00 1.00 1.00 .996 .987 1.64$\pm$.04
CombTopK5_200 .979 .984 .981 .892 .818 20.29$\pm$.05
CombTopK50_500 .993 .989 .994 .975 .961 21.57$\pm$.06
CombTopK100_1000 1.00 .997 .996 .993 .992 23.00$\pm$.08
Table 4: Evaluation of top $K$ retrieval performance, with hit rate (HR) normalized by the brute-force top $K$ method and latency with standard deviation measured over the evaluation of retrieval for a batch of queries (where the batch size is 32). (Relative) hit rate higher than .99 is marked in bold.
We evaluate the following methods on their top $K$ retrieval performance:
  • Brute-force top $K$ (BruteForce): Evaluate the MoL scores for all items and return the top $K$ items. This is the ground truth in our top $K$ evaluation  [^1] We omit the baseline with the two-pass exact algorithm (Section 3.1) because the range-based item retrieval can still be expensive when the range threshold is loose. Empirically, the brute-force top $K$ is more efficient on our datasets. We leave the efficient implementation of the two-pass exact algorithm as future work..
  • Per embedding top $K$ (TopKPerEmbd($N$)): This algorithm is described in Section 3.2. $N$ is the number of candidate items retrieved from each embedding set, where $N\times P\geq K$.
  • Average top $K$ (TopKAvg($N$)): This algorithm is described in Section 3.2. $N$ is the number of the candidate items retrieved by average dot products, where $N\geq K$.
[…]
Table 4 shows the hit rate (HR) and latency of all the methods. The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$. We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023). We set batch size to $32$ for ML-1M, ML-20M, and Books. We perform evaluation on a single NVIDIA RTX 6000 Ada GPU. We report latency averaged over 20 warm runs.
We observe that our approximate heuristics achieve high HR with oversampling. For example, TopKAvg500 is >.$>.99$ in relative HR across the board for ML-1M and ML-20M, and TopKAvg4000 is >.$>.99$ in relative HR across the board for Books. In addition, the combined top $K$ algorithm can outperform both TopKPerEmbd and TopKAvg of the corresponding configurations, sometimes significantly, e.g., CombTopK5_200 vs. TopKPerEmbd5 and TopKAvg200 on Books. This indicates that the set of candidate items retrieved by each individual approximate algorithm indeed complements each other when the weight distributions vary in MoL.
In terms of efficiency, we observe that our approximate heuristics are significantly lower in latency than the exact baseline, especially as the number of items in the dataset becomes large. For example, compared to BruteForce, TopKAvg achieves >.$>.99$ relative HR@100 with a speedup of 4×4$\times$4 × and 91×91$\times$91 × in latency for ML-20M and Books, respectively. While the latency of the retrieval grows with the size of the dataset in the brute-force baseline, it grows much slower with the approximate methods. For example, the latency of retrieval increases by 4.8×4.8$\times$4.8 × from ML-1M to ML-20M in BruteForce, while the growth rate is 1.8×1.8$\times$1.8 × and $\times1$ for TopKPerEmbd100 and TopKAvg1000 respectively. Thus, we expect that the speedup of the approximate methods to become even more prominent as the size of the dataset further increases.
We also notice that TopKAvg tends to be more efficient than TopKPerEmbd with comparable HR, e.g., TopKAvg2000 vs. TopKPerEmbd100 on Books with $\times2$ speedup in latency. We believe that this is mainly due to two reasons. First, when the HR is comparable, the maximal number of candidate items from TopKPerEmbd is larger than that of TopKAvg. Second, compared to TopKPerEmbd, the computation of TopKAvg is agnostic to the number of embedding sets because of the materialization optimization described in Section 3.2. Interestingly, we also see that the combined top $K$ is more efficient than the summation of the latency of its individual components, e.g., CombTopK5_200 is 1.5×1.5$\times$1.5 × faster than the sum of the latency from TopKPerEmbd5 and TopKAvg200 on ML-20M. This is because our implementation reduces the overhead of the combined method by consolidating some shared processing of the two components.
Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ can be used if the target HR is extremely stringent.
Limitations
[…]
5 Related work
Retrieval in Recommendation Models.
Most modern recommendation models follow a classical two-stage information retrieval paradigm (Covington et al., 2016), where up to billions of items (Eksombatchai et al., 2018; Li et al., 2019) are first filtered down to hundreds to thousands in the retrieval or candidate generation stage, followed by the ranking stage that produces the final top 1-10 candidates. Earlier work on scaling large-scale recommendation models primarily utilize dual-encoder setups(Covington et al., 2016) or sometimes graph algorithms (Eksombatchai et al., 2018). Researchers quickly realized that dot products limited retrieval stage’s performance, and introduced multi-query embeddings (Li et al., 2019) and specialized neural networks (He et al., 2017) to alleviate the expressiveness constraints in retrieval. Learned index structures, or generative retrieval, map each candidate item to a series of tokens and leverage beam search to replace approximate nearest neighbor search. These tokens are represented either implicitly during tree traversal (Zhu et al., 2018; Zhuo et al., 2020) or explicitly (Gao et al., 2021). However, they introduce stability challenges as recommendation systems need to support billion-scale dynamic realtime updated set of items. A latest trend is to exploit characteristics of GPUs, TPUs, and other accelerators; learned similarity functions as well as hardware-aware algorithms have been shown to achieve significant quality improvements (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023).
Efficient Nearest Neighbor Search (NNS).
Nearest neighbor search has been a popular topic of research due to their critical role in large-scale retrieval and vector databases. Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS). Various techniques were proposed and analyzed, including tree structures (Bentley, 1975; Ram and Gray, 2012), locality sensitive hashing (Gionis et al., 1999; Shrivastava and Li, 2014), production quantization (Jegou et al., 2011), data partitioning (Li et al., 2002; Zhai et al., 2011), graph-based methods (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019), and so on. The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees while maintaining max-heap property have been proposed to ensure beam search result in globally optimal top-$K$ results (Zhuo et al., 2020). Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques (Shrivastava and Li, 2014; Guo et al., 2016) no longer fully utilize GPUs; to address this issue, accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024).
6 Conclusion
We have analyzed techniques for efficient retrieval with learned similarities in this work. We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and empirically MoL consistently outperforms dot products in HR across 6 scenarios by an average of 21.4% in HR@1 and 13.7% in HR@10, setting new state-of-the-art on recommendation datasets. We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds. Across all datasets evaluated, we demonstrate that our approximate top $K$ algorithms can reach .99.99.99.99 of Hit Rate relative to exact algorithms, while achieving up to 91×91$\times$91 × reduction in end-to-end latency and with minimal indexing overheads. We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations. We believe this result is of interest to both recommendation system research communities and researchers working on enabling advanced learned similarities in vector databases.